OpenStudy (anonymous):

Give examples of functions that each has the following symmetry: a) odd b) even c) none

OpenStudy (anonymous):

@Neemo

OpenStudy (anonymous):

Which one do you think it is?

OpenStudy (anonymous):

odd functions are those f(-x)=-f(x) ex: f(x)=2x f(-x)=2(-x)=-2x=-f(x) so f(x) is odd func even func f(-x)=f(x) ex:f(x)=x^2 f(-x)=(-x)^2=x^2=f(x) f(x) is even func

OpenStudy (anonymous):

a)  b) y= x^2 c) y = x+1

OpenStudy (anonymous):

what about none? @sai1234

OpenStudy (anonymous):

with satisfies both of them is none

OpenStudy (anonymous):

so i dont need towrite anything for c) ?